OpenStudy (anonymous):

Be sure to distribute and collect like terms before you solve. 2(4x + 3) - 9 = 45

OpenStudy (anonymous):

x=6

OpenStudy (mrnood):

@DUDE..LOOK..ITS..SGTC What use is giving "6" as an abswer OP will just have to ask for the NEXT answer and has learned nothing

OpenStudy (anonymous):

sorry, hold on

OpenStudy (mrnood):

@ThebigbossAsb20 the question gives you hints: "distribute" means expand the bracket - multiply the 2 by the elements inside: 2(4x+3) = 2*4x +2*3 write the original equation having expanded as I showed above ...

OpenStudy (anonymous):

First, let's work on the left hand side of your equation, the 2(4x+3)-9 This means, for instance, to see if it can be simplified at all. Multiply x and 4 Multiply x and 1 The x just gets copied along. The answer is x x 4*x evaluates to 4x 4*x+3 evaluates to 4x+3 Multiply 2 by 4x+3 we multiply 2 by each term in 4x+3 term by term. This is the distributive property of multiplication. Multiply 2 and 4x Multiply 1 and x The x just gets copied along. x 2 × 4x = 8x Multiply 2 and 3 1 2 × 3 = 6 2*(4*x+3) evaluates to 8x+6 6 - 9 = -3 The answer is -3+8x 2*(4*x+3)-9 evaluates to -3+8x So, all-in-all, the left hand side of your equation can be written as: -3+8x ________________________________________ Now, let's work on the right hand side of your equation, the 45 The right hand side of your equation can be written as: 45 ________________________________________ So with these (any) simplifications, the equation we'll set out to solve is: -3+8x = 45 Move the -3 to the right hand side by adding 3 to both sides, like this: To the left hand side: -3 + 3 = 0 The answer is 8x To the right hand side: 45 + 3 = 48 The answer is 48 Now, the equation reads: 8x = 48 To isolate the x, we have to divide both sides of the equation by the other variables around the x on the left side of the equation. ________________________________________ The last step is to divide both sides of the equation by 8 like this: To divide x by 1 The x just gets copied along in the numerator. The answer is x 8x ÷ 8 = x 48 ÷ 8 = 6 The solution to your equation is: x = 6

OpenStudy (anonymous):

@ThebigbossAsb20

OpenStudy (anonymous):

@DUDE..LOOK..ITS..SGTC @MrNood I didnt understand it at first. I got negative 6 but i did it again and i got 6 (: thank you guys !

OpenStudy (anonymous):

any time :) , medal please @ThebigbossAsb20